A lamp shade has top and bottom diameters 12cm and 19cm respectively and is 10cm high. Find the height of the cone of which this could be the lower part.
Answer by Cromlix(4381)


Similar triangles
x/(10 + x) = 6/9.5
Cross multiply
9.5x = 6(10 + x)
9.5x = 60 + 6x
3.5x = 60
x = 17.1cm
Therefore height = 10 + 17.1cm
= 27.1 cm